OpenStudy (anonymous):

If you ever swam in a pool and your eyes began to sting and turn red, you felt the effects of an incorrect pH level. pH measures the concentration of hydronium ions and can be modeled by the function p(t) = −log10t. The variable t represents the amount of hydronium ions; p(t) gives the resulting pH level.

OpenStudy (anonymous):

@campbell_st @Michele_Laino

OpenStudy (anonymous):

1. Create a graph of the pH function either by hand or using technology. Locate on your graph where the pH value is 0 and where it is 1. You may need to zoom in on your graph.

OpenStudy (anonymous):

but i still dont understand how to do it:(

OpenStudy (anonymous):

@Michele_Laino @mathway @mathmate

OpenStudy (anonymous):

@Elsa213 Thanks for your help but u should delete this b4 someone tries 2 report u

OpenStudy (anonymous):

Lol okay :) Just trying to help<3

OpenStudy (mathmate):

For the part finding t when p(t)=0.5 we solve the equation p(t)=0.5=-log10(t) raise both sides of the equation to the base 10, \(10^{0.5} = 10^{-log10(t)}= 10^{log10(1/t)}=1/t\) gives \(\sqrt{10}=1/t\) or \(t=1/\sqrt{10}=\sqrt{10}/10\) check: \(p(\sqrt{10}/10)=-log10(sqrt{10}/10)=\)=1/2